36-303: Sampling, Surveys & Society HW04: Hand in March 3, BY EMAIL ONLY

Reminders/Announcements:

No Class Thu Mar 3 BJ will be out of town. Note that there are things to turn in (by email!) on Mar 3.

Team project assignments (see updated team project assignment sheet, on class website):

• II.5a Sampling Scheme & Question Design (Thu Feb 24): Email items K, L, M to brian@stat.cmu.edu; more details on revised team project sheet, and in lecture 13.

• II.5b Pretest & Revise Questions (Tue Mar 1): Email items N, O to brian@stat.cmu.edu

• II.6 Final IRB & Project Plan (Thu Mar 3): Email the items listed on the revised team assignment sheets to brian@stat.cmu.edu

Reading:

• Read/Review all of Groves Ch 4: Sample Design & Sampling Error

• Read/Review Groves Ch 6: Nonresponse

• New: Groves Ch 10: Post-Survey Processing

• If you have Lohr, you could see some gory details in Ch 2 (SRS), Ch 4 (Stratified Sampling) and Ch 5 (Cluster Sampling)

Exercises to Turn In (there is only 1 exercise):

1. Peer evaluations.

(a) Download and make several copies of the peer evaluation sheet from the class website [pdf or doc]. If you are in a team of 4 students, you will need 3 copies. If you are in a team of 5 students, you will need 4 copies, etc.

(b) Fill out the peer evaluation electronically (using msword or a pdf editor) for each of your team-mates in this class.

(c) Send a single email to brian@stat.cmu.edu with

• Title: 36-303 peer evaluation

• All the peer evaluations that you filled out all attached to it.

You will be graded both on whether you do a good job of filling out your peer evaluations, and on the evaluation of you by others on your tesm.